OpenStudy (anonymous):

A pharmacist wants to make 20 ml of a 59% hydrogen peroxide solution. How many mls of 35% hydrogen peroxide must he mix with 75% hydrogen peroxide to make the desired solution?

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

the pharmacist needs how many of 35%???

OpenStudy (anonymous):

and 75%???

OpenStudy (anonymous):

@AntonioCochrane

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

@FriedRice

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Let the mixture be made from "x" ml of 35% hydrogen peroxide and (20-x) ml of 75% hydrogen peroxide. x ml of 35% hydrogen peroxide will have 0.35x of hydrogen peroxide. (20-x) ml of 75% hydrogen peroxide will have 0.75(20-x) of hydrogen peroxide. Total hydrogen peroxide = 0.35x + 0.75(20-x) = 0.35x + 15 - 0.75x = 15 - 0.4x The mixture has (15 - 0.4x) ml of hydrogen peroxide in 20 ml of solution. The concentration of the mixture is 59%. Therefore, 15 - 0.4x = 20 * .59 = 11.8 15 - 11.8 = 0.4x 0.4x = 3.2 x = 8 So the pharmacist should mix 8 ml of 35% hydrogen peroxide with 12 ml of 75% hydrogen peroxide to get the 20 ml of a 59% hydrogen peroxide solution.